OpenStudy (anonymous):

Systems of Linear Equations. Add/Elimination. Line up variables and solve for the other. Give Solution and show your work. x+3y=13 & x+y=5

OpenStudy (unklerhaukus):

\[x+3y=13\qquad_{(i)}\]\[x+y=5\qquad_{(ii)}\] \[x+3y-x+y=13-5\qquad_{(i)-(ii)}\]

OpenStudy (anonymous):

the way I would do it is to first solve for y. So x+y = 5 y = 5-x Plug that into the other equation and let me know what you get.

OpenStudy (anonymous):

OH its elimination by addition - two seconds :) im sorry about that!

OpenStudy (anonymous):

but meanwhile if you'd like you can find the answer via substitution so that as we go through the elimination process you can see why we got the answer and how

OpenStudy (anonymous):

I just figured it out. I would have to multiply equation 2 by -3 in order to add or subtract the variables. I found my answer .... I used equation to double check and it worked. I got (1,4) for x,y

OpenStudy (anonymous):

oh awesome! :) good job!

OpenStudy (anonymous):

thanks